OpenStudy (anonymous):

A rectangle's sides are measured to be 6.2 cm and 9.3 cm. What is the rectangle's area rounded to the correct number of significant digits? a. 57.66 cm squared b. 57.7 cm squared c. 58 cm squared d. 60 cm squared

OpenStudy (anonymous):

c. 58 cm squared

OpenStudy (anonymous):

What exactly does the last part mean? 'rounded to the correct number of significant digits'

OpenStudy (anonymous):

a=6.2(2significant digits)*9.3(2significant digits)=58(2significant digits)cm^2

OpenStudy (anonymous):

@ASAAD123 Thank you so much.